If the value of K_c for the reaction : is 7.5 Tim
Answer #1

9) Kc = [NO]2/[N2][O2] = 7.5 x 10-9

Kc' = [N2][O2]/[NO]2 = 1/Kc =1/ 7.5 x 10-9 = 1.33 x 108

The answer is 7.5 x 10-9.

10) i) Increase the concentration of H2.

ii) Increase the concentration of N2.

iii) Decrease the concentration of NH3.

iv) Decrease the temperature of the reaction.

v) Increase the pressure of the reaction.